OpenStudy (anonymous):

The radius of the cone is 1.25 inches, and its height is 2.75 inches. If the diameter of the bubble gum ball is 0.5 inches, what is the closest approximation of the volume of the cone that can be filled with flavored ice? 4.43 in3 0.07 in3 13.50 in3 4.50 in3

OpenStudy (valpey):

Volume of Cone \[\frac{1}{3}\pi r^2h\] Volume of Sphere \[\frac{4}{3}\pi r^3\] Volume of Ice = Volume of Cone - Volume of Sphere